(D-1)2(D2+1)2 y= sin 2(x/2)+ex+x

1
Expert's answer
2021-06-09T09:18:14-0400

"y(x)=y_c+y_p"

General solution:

"y_c=c_1e^x+c_2xe^x+c_3cosx+c_4sinx+x(c_5cosx+c_6sinx)"


For particular solution:

"sin^2(x\/2)=(1-cosx)\/2"

Then:


Particular solution:

"PI=\frac{1}{f(D)}(sin^ 2(x/2)+e^x+x)"

"PI_1=\frac{1}{f(D)}e^x=\frac{x^2e^x}{2!(1+1)^2}=\frac{x^2e^x}{8}"


"PI_2=\frac{1}{f(D)}(-\frac{1}{2}cosx)=-\frac{1}{2}(\frac{1}{-2D}cosx+\frac{x}{4}\int xsinxdx)="


"=\frac{sinx}{4}-\frac{x(sinx-xcosx)}{4}"


"PI_3=\frac{1}{f(D)}(x+1/2)"


"(D-1)^{-2}=1+2D+..."

"(D^2+1)^{-2}=1-2D+..."


"PI_3=(1-2D)(x+2.5)=x+2.5-2=x+1\/2"


"y_p=PI_1+PI_2+PI_3"